OpenStudy (anonymous):

d/dx int (u^2-1)/(u^2+1) from 2x to 3x

OpenStudy (anonymous):

\[d/dx \int\limits (u^2-1)/(u^2+1) from 2x \to 3x\]

OpenStudy (amistre64):

... i forgot something lol

OpenStudy (amistre64):

\[\frac{d}{dx}\left(\int_{2x}^{3x}(\frac{u^2-1}{u^2+1})\right)\]

OpenStudy (anonymous):

Yep

OpenStudy (amistre64):

well, there is a way to use the fundamental thrm to this, but I never pay attention to it, makes life way to easy :)

OpenStudy (anonymous):

I end up going through the calculations and getting \[(3* (3x-1)/(3x^2+1)) + (2* (2x^2-1)/(2x^2+1)) \]

OpenStudy (anonymous):

3x^2*

OpenStudy (amistre64):

it ints up to: \[F(3x)-F(2x)\] and then deriving back down to this little equation

OpenStudy (amistre64):

the F(x) is really inconsequantial to the process; but I cant recall how

OpenStudy (amistre64):

\[\int_{2x}^{3x}f'(u)du \implies f(3x)-f(2x)\] \[\frac{d}{dx}(f(3x)-f(2x)) \implies f'(3x)*3 - f'(2x)*2\] \[f'(u)=\frac{u^2-1}{u^2+1}\] \[\frac{(3x^2)-1}{(3x^2)+1}-\frac{(2x^2)-1}{(2x^2)+1}\] right?

OpenStudy (amistre64):

ack... all that work and I left out the *3 and *2 lol

OpenStudy (anonymous):

The answer turns out be [according to the solution manual [(-2 * 4x^2-1/4x^2+1)+(3* 9x^2-1/9x^2+1)\]

OpenStudy (amistre64):

\[\frac{3((3x)^2-1)}{(3x)^2+1}-\frac{2((2x)^2-1)}{(2x)^2+1}\]

OpenStudy (amistre64):

yeah, when I get all the parts right; I see it happening :)

OpenStudy (anonymous):

My question is where does the second 3 and 2 come from that you multiply.

OpenStudy (amistre64):

the chain rule really....

OpenStudy (anonymous):

but, the 2 and 3 are already in the equation and aren't being multiplied..

OpenStudy (amistre64):

it doesnt really matter what the function integrates up to; because you are simply bringing it back down to the derivative again right?

OpenStudy (anonymous):

Yea, it's just putting it into terms of x .

OpenStudy (anonymous):

Oh snap, never mind I see it.

OpenStudy (amistre64):

\[F(x)=\int_{2x}^{3x}\left(\frac{u^2-1}{u^2+1}\right)du\] and then you work in the \(F(3x)-F(2x)\) and derive it back again

OpenStudy (anonymous):

Herp derp, thanks a bunch.

OpenStudy (amistre64):

:) youre welcome